OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)? x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

OpenStudy (decentnabeel):

standard form of circle: (x - h)^2 + (y - k)^2 = r^2 center is (h , k) and radius is r for your circle, the center (h , k) is (-2 , 1) so it will look like: (x +2)^2 + (y - 1)^2 = r^2 to find the radius squared, plug in the (x,y) point you know is on the circle -- (-4,1) (-4 +2)^2 + (1 - 1)^2 = r^2 -2^2 + 0 = 4 = r^2 thus, the equation is: (x +2)^2 + (y - 1)^2 = 4

OpenStudy (decentnabeel):

x^2+4+4x+y^2+1-2y=4 x^2+y^2+4x-2y+1=0 so second opton is the answer

OpenStudy (decentnabeel):

are you understand @king12233